Please help me answer this question, appreciate it a lot.

1.A 1Kb deletion in geneA was found to be responsible for a human recessive genetic disease. A diagnostic PCR reaction was designed to amplify part of gene A (3Kb) flanking the deletion site. Please provide predicted PCR results for non-carrier and recessive disease.

b. What happens to depth of field as you increase numerical aperture?

c. To make a liter 10 X TBE buffer, you need 890 mM Tris (MW 121g/mol), how many grams of Tris do you need to weight out?

A. 108 g Tris B. 55gTris C. 9.3 Tris D. 10 g Tris

d. To make 30 mL of 0.7% agarose gel, how many grams of agarose do you need to weight out?

A. 0.7 g B. 0.2 g C. 0.3 g D. 1 g
Answer #1

Answer d) 0.7 % means 0.7 g in 100 ml

So if 100 ml have 0.7 g

30 ml have = 0.7 ×30 / 100 = 0.21 g

So option B is correct

Answer c) 1 mol of tris = 121 g

1 millimol of tris = 121/1000 = 0.121 g

890 millimol of tris = 0.121 ×890 = 107.670

= 108 g